# Reinforcement Learning with Action Sequence for Data-Efficient Robot Learning

### v1

@@ 6 Conclusion
We presented Coarse-to-fine Q-Network with Action Sequence (CQN-AS), a value-based RL algorithm that trains a critic network that outputs Q-values over action sequences. Extensive experiments in benchmarks with various setups show that our idea not only improves the performance of the base algorithm but also allows for solving complex tasks where prior RL algorithms completely fail.
We believe our work will be strong evidence that shows RL can realize its promise to develop robots that can continually improve through online trial-and-error experiences, surpassing the performance of BC approaches. We are excited about future directions, including real-world RL with humanoid robots, incorporating advanced critic architectures (Kapturowski et al., 2023; Chebotar et al., 2023; Springenberg et al., 2024), bootstrapping RL agents from imitation learning (Hu et al., 2023; Xing et al., 2024) or offline RL (Nair et al., 2020; Lee et al., 2021), extending the idea to recent model-based RL approaches (Hafner et al., 2023; Hansen et al., 2024), to name but a few.
@@ Acknowledgements
We thank Stephen James and Richie Lo for the discussion on the initial idea of this project. Pieter Abbeel holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon.
@@ Appendix A Experimental Details
We use the same set of hyperparameters across the tasks in each domain. For hyperparameters shared across CQN and CQN-AS, we use the same hyperparameters for both algorithms for a fair comparison. We provide detailed hyperparameters for BiGym and RLBench experiments in Table 3 and hyperparameters for HumanoidBench experiments in Table 4
@@
Batch size 128 (BiGym) / 256 (RLBench)
Demo batch size 128 (BiGym) / 256 (RLBench)
@@
Dueling network False
@@
Target critic update interval ($\tau$) 1000
@@
For all experiments, we use consumer-grade 11GB GPUs such as NVIDIA GTX 1080Ti, NVIDIA Titan XP, and NVIDIA RTX 2080Ti with 11 or 12GB VRAM. With 2080Ti GPU, each BiGym experiment with 100K environment steps take 9.5 hours, each RLBench experiment with 30K environment steps take 6.5 hours, and each HumanoidBench experiment with 7M environment steps take 48 hours. We find that CQN-AS is around 33% slower than running CQN because larger architecture slows down both training and inference.

### v2

@@ 6 Conclusion
In this paper, we presented Coarse-to-fine Q-Network with Action Sequence (CQN-AS), a value-based RL algorithm that trains a critic network that outputs Q-values over action sequences. Extensive experiments in benchmarks with various setups show that our idea not only improves the performance of the base algorithm but also allows for solving complex tasks where prior RL algorithms completely fail.
We believe our work will be strong evidence that shows RL can realize its promise to develop robots that can continually improve through online trial-and-error experiences, surpassing the performance of BC approaches. We are excited about future directions, including real-world RL with humanoid robots, incorporating advanced critic architectures (Kapturowski et al., 2023; Chebotar et al., 2023; Springenberg et al., 2024), bootstrapping RL agents from imitation learning (Hu et al., 2023; Xing et al., 2024) or offline RL (Nair et al., 2020; Lee et al., 2021), extending the idea to recent model-based RL approaches (Hafner et al., 2023; Hansen et al., 2024), fine-tuning vision-language-action models that use action sequence (Team et al., 2024; Doshi et al., 2024; Kim et al., 2024) with our algorithm, to name but a few.
@@ Acknowledgements
We thank Stephen James and Richie Lo for the discussion on the initial idea of this project. This work was supported in part by Multidisciplinary University Research Initiative (MURI) award by the Army Research Office (ARO) grant No. W911NF-23-1-0277, an ONR DURIP grant, and BAIR Industrial Consortium. Pieter Abbeel holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon.
@@ Appendix A Experimental Details
We use the same set of hyperparameters across the tasks in each domain. For hyperparameters shared across CQN and CQN-AS, we use the same hyperparameters for both algorithms for a fair comparison. We provide detailed hyperparameters for BiGym and RLBench experiments in Table 3 and HumanoidBench experiments in Table 4
@@
Batch size 256
Demo batch size 256
@@
Dueling network True
@@
Target critic update interval ($\tau$) 100
Update-to-data ratio (UTD) 0.5
@@
For BiGym and Humanoid experiments, we use NVIDIA A5000 GPU with 24GB VRAM. With A5000, each BiGym experiment with 100K environment steps take 16 hours, and each HumanoidBench experiment with 10M environment steps take 40 hours. For RLBench experiments, we use NVIDIA RTX 2080Ti GPU, with which each experiment with 30K environment steps take 6.5 hours. We find that CQN-AS is around 33% slower than running CQN because larger architecture slows down both training and inference.